a playing field is a rectangle that is 800 feet long by 180 feet wide. Find, to the nearest foot, the length of a straight-line run that started at one corner and went diagonally to end at the opposite corner
Answer by fcabanski(1391)

The diagonal is the hypotenuse of a right triangle with the legs the width and length of the rectangle.


Find the hypotenuse using the Pythagorean Formula: Hypotenuse squared = the sum of the squared sides: a^2 + b^2 = c^2.


diagonal^2 = 800^2 + 180^2 = 640000 + 32400 = 672400.


Diagonal (hypotenuse) is the square root of 672400 = 820 feet.